### v3

# Adler function, Bjorken polarized Sum Rule: confirmation of elements of the $\{\beta\}$-expansion

Abstract
Different ways exist to obtain the elements of the $\{\beta\}$-expansion for renormgroup invariant quantities. Here we consider independent confirmation within the standard QCD of a number of our results Baikov:2022zvq for the values of elements of this expansion for the nonsinglet Adler $D_{A}$-function, Bjorken polarized sum rules $S^{Bjp}$ up to the order N4LO. We suggest an approach to estimate the results of high order QCD calculations using a smaller number of diagrams of the specific type. This type is based on a proposed generalization of Naive NonAbelianization.
Keywords:
Renormalization Group, QCD
October 30, 2024
1 Introduction
Here we will consider the results of $\{\beta\}$–expansion Mikhailov:2004iq and discuss the reliability of the obtained elements of this expansion Baikov:2022zvq and the relation of the latter with the corresponding Feynman diagrams. The elements of this expansion appear at the decomposition of the standard perturbative coefficients of any renormalization group invariants (RGIs); here these coefficients are $d_{n}$ or $c_{n}$ for the nonsinglet Adler $D_{A}$-function, or the Bjorken polarized sum rule (BpSR) $S^{Bjp}$ respectively. We consider independent confirmation of our results for the decomposition of the nonsinglet Adler $D_{A}$-function for the values of elements $d_{3}[.]$ ($c_{3}[.]$) of $\{\beta\}$–expansion and for some elements of $d_{4}[.]$ Baikov:2022zvq . The agreements are obtained using the already known results of N3,4LO calculations Ball:1995ni of some diagrams in the $\overline{\rm MS}$scheme in pure QCD. In the case of the nonsinglet $D_{A}$-function at the default choice of scale $Q^{2}=\mu^{2}$,
$D_{A}(a(\mu^{2}))=1+\sum_{n\geq 1}a^{n}_{s}(\mu^{2})~{}d_{n},~{}~{}a_{s}(\mu^{2})\!=\!\alpha_{s}(\mu^{2})/(4\pi),$ (1)
the decomposition of $d_{n}$ looks like Mikhailov:2004iq ; Baikov:2022zvq ,
$\vspace*{-2mm}a_{s}^{2}~{}~{}d_{2}$ $=$ $\!\beta_{0}\,d_{2}[1]+d_{2}[0]\,,$ (2a)
$a_{s}^{3}~{}~{}d_{3}$ $=$ $\!\beta_{0}^{2}\,d_{3}[2]+\beta_{1}\,d_{3}[0,1]+\beta_{0}\,d_{3}[1]+d_{3}[0]\,,$ (2b)
$a_{s}^{4}~{}~{}d_{4}$ $=$ $\!\beta_{0}^{3}\,d_{4}[3]+\beta_{1}\,\beta_{0}\,d_{4}[1,1]+\beta_{2}\,d_{4}[0,0,1]+\beta_{0}^{2}\,d_{4}[2]+\beta_{1}d_{4}[0,1]+\beta_{0}\,d_{4}[1]$ (2c)
[…]
$\ldots$ $=$ $\ldots$
$a_{s}^{n}~{}~{}d_{n}$ $=$ $\beta_{0}^{n-1}d_{n}[n-1]+\ldots+d_{n}[0]\,,$ (2d)
where $\beta_{i}$ are the coefficients of expansion of the QCD $\beta$-function (see Appendix A). The $\{\beta\}$-expansion shows how the intrinsic charge renormalization manifests itself; thus all possible $\beta_{i}$-terms appear in each order $n$ of perturbative expansion (PT). This property differ it from more or less formal expansions over powers of selected lagrangian parameters or Casimirs, say the number of active quark flavors $n_{f}$, or powers of $\beta_{0}$ Lovett-Turner:1995zwc , which is based on the same formal reasons. The procedure of calculating the expansion elements in (2) was based, in its turn, on the results of extended QCD (QCDe) presented in Chetyrkin:2022fqk , which possesses a number of new degrees of freedom (d.o.f.). These d.o.f. are arbitrary number of different fermion representations, revealing themselves only in intrinsic loops. After decomposition as in (2), one can equate the contributions of these new d.o.f. to zero and return from QCDe to the standard QCD. Note here that now are at least two different approaches to obtain the values of the elements within the standard QCD that lead to different results. One of them, started with Mojaza:2012mf , is based on an interpretation of RG transformation, it got the name “Principle of Maximum Conformality” (PMC), see review DiGiustino:2023jiq and references therein. The other approach, see Cvetic:2016rot , is based on an extended interpretation of Crewther-Broadhurst-Kataev (CBK) relation Broadhurst:1993ru . So, some doubt appeared about how reliable the results for the elements obtained within QCDe in Baikov:2022zvq are with respect to the standard QCD.
Here we confirm our results in order $O(a_{s}^{3})$ and confirm them in part in order of $O(a_{s}^{4})$ based on the calculation within the standard QCD and focusing on the diagrammatic origins of the elements of the $\{\beta\}$-expansion. In addition in Sec.4 we propose a hierarchy of contributions to the coefficients $d_{n}(c_{n})$ based on a certain generalization of Naive Nonabelianization (NNA) in the framework of the $\{\beta\}$-expansion. The usage of this hierarchy can facilitate the laborious process of estimating physical quantities in high-loops calculations.
2 Test for $\{\beta\}$-expansion elements of $D_{A}\,(S^{Bjp})$ at $O(a_{s}^{3})$
Recall the elements of the $\{\beta\}$-expansion for $D_{A}$ up to the order $O(a_{s}^{3})$ Mikhailov:2004iq ; Baikov:2022zvq ; Kataev:2014jba , where the first term $d_{1}=3\,C_{\text{F}}$ :
$d_{2}[1]$ $=$ $d_{1}\left(\frac{11}{2}-4\zeta_{3}\right);~{}d_{2}[0]=d_{1}\left(\frac{\rm C_{A}}{3}-\frac{\rm C_{F}}{2}\right);$ (3)
$d_{3}[2]$ $=$ $d_{1}\left(\frac{302}{9}-\frac{76}{3}\zeta_{3}\right);~{}d_{3}[0,1]=d_{1}\left(\frac{101}{12}-8\zeta_{3}\right);$ (4a)
$d_{3}[1]$ $=$ $d_{1}\left[{\rm C_{A}}\left(-\frac{3}{4}+\frac{80}{3}\zeta_{3}-\frac{40}{3}\zeta_{5}\right)-{\rm C_{F}}\left(18+52\zeta_{3}-80\zeta_{5}\right)\right];$ (4b)
$d_{3}[0]$ $=$ $d_{1}\Bigg{[}{\rm C_{A}^{2}}\left(\frac{523}{36}-72\zeta_{3}\right)+\frac{71}{3}{\rm C_{A}C_{F}}-\frac{23}{2}{\rm C_{F}^{2}}\Bigg{]}~{}.$ (4c)
These decompositions for $d_{3}$ were first obtained in Mikhailov:2004iq where one used the results for $D_{A}$ with a single additional degree of freedom – the light gluino Chetyrkin:1996ez , then the elements of this decomposition were confirmed in Baikov:2022zvq within QCDe with any numbers of fermion d.o.f. Below we reproduce the same values of the elements $d_{3}[.]$ by considering the results for specific diagrams for $D_{A}$ calculated in pure QCD.
Let us start with the simplest case in (3). The first element $d_{2}[1]$ originates from the diagram where the single gluon line is dressed in a quark bubble, which is proportional to $T_{R}n_{f}\sim\beta_{0}$. Just this correspondence allows one to perform the decomposition unambiguously. The sequential dressing of this only one gluon line generates its anomalous dimension that is finally revealed as a set of $\beta$-elements of our interest also in the higher orders of PT.
Refer to caption           Refer to caption
Figure 1: Diagrams for $D_{A}$ in order $a_{s}^{3}$ with self-energy dressing of only one gluon line (red in color). One should add to these samples of diagrams all possible connections of gluon lines. Left. proportional to $C_{\text{F}}\cdot(T_{R}n_{f}\equiv x)^{2}$, $\sim\beta_{0}^{2}$. Right. proportional to $C^{2}_{\text{F}}(T_{R}n_{f}\equiv x)$, $\sim\beta_{1}$  .
In the next order the 2nd loop dressing of only one gluon line generates specific diagrams with two 1-loop bubbles (reducible), see Fig.1(Left) and with 2-loop one (irreducible), Fig.1(Right).
We reproduced in Fig.1 the diagrams taken from Fig.C1(IV1) and Fig.C1(IV2) in Ball:1995ni , and the abelian projection of the results corresponding to them was presented in Appendix C, Eqs.(C.2),(C.3) there. The result for the diagram in Fig.1(Left) independently confirms the value of the $d_{3}[2]$ term in (4a) and coincides also with the general expression for $d_{n}[n-1]$, which can be extracted from the results of Broadhurst:1992si ; Broadhurst:1993ru . The diagram in the right panel contributes to the term $\beta_{1}d_{3}[0,1]$ by means of anomalous dimension of a gluon in order of $a_{s}^{2}$. The latter, in its turn, includes the unique Casimir $C_{\text{F}}\cdot(T_{R}n_{f}\equiv x)$, which is common with $\beta_{1}$, see Eq.(24b). This diagrammatic correspondence unambiguously determines the element $d_{3}[0,1]$ within the standard QCD calculation that value coincides with the one in (4a). Further we will trace the contributions to $d_{n}$ of the color structures starting with maximum numbers of quark loops: the evident case of a renormalon chain with only one gluon line contribution gives us $C_{F}\cdot x^{(n-1)}$, then one fermion loop less – $C_{F}(C_{F}x^{(n-2)})$, then $C_{F}(C_{F}^{2}x^{(n-3)})$ and so on. In other words, these terms are a projection onto the abelian sector of the QCD calculation (proportional to QED) of $D_{A}$ and $S^{Bjp}$. The abelian part of the structure of $\beta$-coefficients is well recognized in the first terms (underlined) of the corresponding formulas in Appendix A.
Now we will show that independent knowledge of $d_{3}[0,1]$ guarantees also independent determination of two remainder elements $d_{3}[1],d_{3}[0]$ of this order if the coefficient $d_{3}$ as a whole is already known. Using the definition $T_{R}n_{f}=x$, we introduce Mikhailov:2016feh the consequence of roots $x_{0},x_{1},\ldots\sim x$ of the QCD $\beta$-function coefficients $\beta_{0}(x_{0})\!=\!0$, $\beta_{1}(x_{1})\!=\!0,\ldots$, respectively,
$\!\!\!\!\!\!\!\!x_{0}\!=\!\displaystyle\frac{11}{4}{\rm C_{A}};$ (5a)
$\!\!\!\!\!\!\!\!\beta_{0}(x_{0})=0;$ $\!\beta_{1}(x_{0})\!=\!-{\rm C_{A}}\!\left(7{\rm C_{A}}+11{\rm C_{F}}\right)\!;$ $\!\!\beta_{2}(x_{0})\!=\!{\rm C_{A}}\!\!\left(\!\frac{11}{2}{\rm C_{F}^{2}-\frac{1127}{24}{\rm C_{A}^{2}}\!-\!\frac{77}{3}{\rm C_{A}}{\rm C_{F}}\!}\!\right)\!\!;$ (5b)
$\displaystyle x_{1}\!=\!\frac{17}{2}\frac{\rm C^{2}_{A}}{5{\rm C_{A}}+3{\rm C_{F}}};$ (5c)
$\!\!\!\!\!\!\!\!\beta_{0}(x_{1})\!=$ $\!\!\!\displaystyle{\rm C_{A}}\frac{7{\rm C_{A}}+11{\rm C_{F}}}{5{\rm C_{A}}+3{\rm C_{F}}};\beta_{1}(x_{1})=0;$ $\beta_{2}(x_{1})=\ldots\,.$ (5d)
Introducing a reduced $\tilde{d}_{3}(x)\!\equiv\!d_{3}(x)-\beta_{0}^{2}(x)d_{3}[2]$, depending on the argument $x$, we obtain from the $\{\beta\}$-expansion for $d_{3}$ (Eq.(1.3) in Baikov:2022zvq ),
$\tilde{d}_{3}(x_{0})-\beta_{1}(x_{0})d_{3}[0,1]$ $=$ $d_{3}[0];$ (6a)
$\frac{\tilde{d}_{3}(x_{0})-\tilde{d}_{3}(x_{1})}{\beta_{0}(x_{1})}+\frac{\beta_{1}(x_{0})}{\beta_{0}(x_{1})}d_{3}[0,1]$ $=$ $d_{3}[1]\,,$ (6b)
where $d_{3}(x)$ (and $d_{4}(x)$) are known from the original calculations Baikov:2010je . So we get independent confirmation within QCD of all the results for $d_{3}[.]$ displayed in Eqs.(4).
It is clear that the same relations, like Eq.(6), is true also for the elements $c_{3}[.]$ in the $\{\beta\}$-expansion of $S^{Bjp}$. The required value of $c_{3}[0,1]$ can be obtained from $d_{3}[0,1]$ based on the CBK relation Kataev:2010du ; Baikov:2022zvq ,
$\!c_{2}[1]+d_{2}[1]\!$ $=$ $\!c_{3}[0,1]+d_{3}[0,1]\!=\!\!c_{n}[\underbrace{0,0,\ldots,1}_{n-1}]+d_{n}[\underbrace{0,0,\ldots,1}_{n-1}]\!=\!3C_{\text{F}}\!\left(\frac{7}{2}-4\zeta_{3}\right).$ (7)
Therefore, the knowledge of this $d_{3}[0,1]$ makes it possible to restore both $d_{3}[.]$ and $c_{3}[.]$ only within QCD calculations and confirms the results obtained with the help of QCDe in Baikov:2022zvq , and within QCD with light gluino Mikhailov:2004iq .
3 Test for $\{\beta\}$-expansion elements of $D_{A}$ at $O(a_{s}^{4})$
Here we discuss independent tests to verify the elements of $\{\beta\}$-expansion for $d_{4}(c_{4})$ obtained in Baikov:2022zvq on the basis of the results presented in Chetyrkin:2022fqk within QCDe. These results for the $d_{4}[.]$ elements are presented below,
$d_{4}[3]$ $=$ $C_{\text{F}}\left(\frac{6131}{9}-406\zeta_{3}-180\zeta_{5}\right);$ (8a)
$d_{4}[1,1]$ $=$ $C_{\text{F}}\left(385-\frac{1940}{3}\zeta_{3}+144\zeta_{3}^{2}+220\zeta_{5}\right);$ (8b)
$d_{4}[0,0,1]$ $=$ $C_{\text{F}}\left(\frac{355}{6}+136\zeta_{3}-240\zeta_{5}\right);$ (8c)
$d_{4}[2]$ $=$ $-C_{\text{F}}\left[C_{\text{F}}\left(\frac{6733}{8}+1920\zeta_{3}-3000\zeta_{5}\right)+\right.$ (8d)
$C_{\text{A}}\left(\frac{20929}{144}-\frac{12151}{6}\zeta_{3}+792\zeta_{3}^{2}+1050\zeta_{5}\right)\right];$
$d_{4}[1]$ $=$ $C_{\text{F}}\bigg{[}-C_{\text{F}}^{2}\left(\frac{447}{2}-42\zeta_{3}-4920\zeta_{5}+5040\zeta_{7}\right)+$ (8e)
$C_{\text{A}}C_{\text{F}}\left(\frac{3301}{4}-678\zeta_{3}-2280\zeta_{5}+2520\zeta_{7}\right)+$
$C_{\text{A}}^{2}\left(\frac{16373}{36}-\frac{17513}{3}\zeta_{3}+2592\zeta_{3}^{2}+3030\zeta_{5}-420\zeta_{7}\right)\bigg{]};$
$d_{4}[0,1]$ $=$ $-C_{\text{F}}\left[C_{\text{A}}\left(\frac{139}{12}+\frac{1054}{3}\zeta_{3}-460\zeta_{5}\right)+C_{\text{F}}\left(\frac{251}{4}+144\zeta_{3}-240\zeta_{5}\right)\right];$ (8f)
$d_{4}[0]$ $=$ $\tilde{d}_{4}[0]+\delta d_{4}$ (8g)
$=$ $C_{\text{F}}^{4}\left(\frac{4157}{8}+96\zeta_{3}\right)-C_{\text{A}}C_{\text{F}}^{3}\left(\frac{2409}{2}+432\zeta_{3}\right)+C_{\text{A}}^{2}C_{\text{F}}^{2}\left(\frac{3105}{4}+648\zeta_{3}\right)+$
$C_{\text{A}}^{3}C_{\text{F}}\left(\frac{68047}{48}+\frac{8113}{2}\zeta_{3}-7110\zeta_{5}\right)+\delta d_{4}\,;$
$\delta d_{4}$ $=$ $-16\Big{[}n_{f}\frac{d_{F}^{abcd}d_{F}^{abcd}}{d_{F}}\left(13+16\zeta_{3}-40\zeta_{5}\right)+\frac{d_{A}^{abcd}d_{F}^{abcd}}{d_{F}}\left(-3+4\zeta_{3}+20\zeta_{5}\right)\Big{]}\,.$ (8h)
Equation(8a) for $d_{4}[3]$ can also be independently confirmed by the partial result for $\beta_{0}^{3}d_{4}[3]$ in Eq.(C.4) in Ball:1995ni , see Fig.C1(V1) there, which agrees with the results in Broadhurst:1992si ; Broadhurst:1993ru too. Here we reproduce these “abelian” diagrams in Fig.2(see left panel) for reader’s convenience.
(i)The projection of the weighted sum $\beta_{2}\,d_{4}[0,0,1]+\beta_{1}\beta_{0}\,d_{4}[1,1]$ onto the Casimir $C_{F}^{2}x^{2}$ is denote as $s_{4}=\hat{P}_{C^{2}_{F}x^{2}}\left(\beta_{2}\,d_{4}[0,0,1]+\beta_{1}\beta_{0}\,d_{4}[1,1]\right)$, where $\hat{P}_{C^{2}_{F}x^{2}}$ means the corresponding projector. The evident diagrammatic origins just of these terms in $s_{4}$ are presented in Fig.C1(V2,V3) Ball:1995ni that is reproduced in Fig.2(Right) here, the corresponding partial results are presented in Eqs.(C5, C6) Ball:1995ni . The results for each of these two diagrams, evidently related to the factors $\beta_{2},~{}\beta_{1}\beta_{0}$ due to gluon line renormalizations, contributes to both the terms in the sum $s_{4}$.
Refer to caption              Refer to caption        Refer to caption
Figure 2: The types of the “abelian” diagrams for $D_{A}$ in order $a_{s}^{4}$ of the self-energy type of only one gluon line (red in color) also include the symmetric diagram where the inner loops are interchanged. Left: proportional to $C_{\text{F}}x^{3}$, $\sim\beta_{0}^{3}$ Right: proportional to $C_{\text{F}}^{2}x^{2}$: $\Rightarrow\beta_{2},\beta_{1}\beta_{0}$
The sum of the partial results of the direct QCD calculation of these diagrams reads in the lhs111 up to the factor $4^{-4}$ in the lhs due to another definition of $a_{s}$, $a_{s}=\alpha_{s}/\pi$ in Ball:1995ni of the equation as
[…]
The abelian projections of the $\beta$-coefficients in the last equation correspond to the underlined terms in Eq.(24) (Appendix A). Substituting the corresponding elements $d_{4}[.]$ from Eqs.(8b,8c) into the rhs, one obtains the equality with the lhs; at the same time, the terms with $\zeta_{5}$ that exist in each of the elements $d_{4}[0,0,1],d_{4}[1,1]$ cancel in the sum $s_{4}$. So the sum of these elements is confirmed. Recall that $\beta_{2}\,d_{4}[0,0,1]+\beta_{1}\beta_{0}\,d_{4}[1,1]$ corresponds to special kinds of dressing of only one gluon line in Fig. 2, but it does not exhaust all the contributions to $C_{F}^{2}x^{2}$.
(ii) To obtain the completed projection on $C_{F}^{2}x^{2}$, presented in the lhs of (10), see Baikov:2010je ,
$\!\!\!\!\!\!\!\!C_{F}^{2}x^{2}4\!\left(\!\frac{5713}{27}-\frac{4648}{3}\zeta_{3}+192\zeta_{3}^{2}+\frac{4000}{3}\zeta_{5}\!\!\right)\!=\!\hat{P}_{C^{2}_{F}x^{2}}\!\Big{(}\!\!\beta_{2}d_{4}[0,0,1]+\beta_{1}\beta_{0}d_{4}[1,1]+\beta_{0}^{2}d_{4}[2]\!\Big{)},$ (10)
one should supplement $s_{4}$ with the contribution $\hat{P}_{C^{2}_{F}x^{2}}\left(\beta_{0}^{2}d_{4}[2]\right)$ in the rhs. The diagrams forming this additional term are not of “the only one dressed gluon line” type. Equality (10) unambiguously determines the abelian part of the element $d_{4}[2]$ and confirms now Eq.(8d). Concluding we confirm the values of the elements $d_{4}[0,0,1],d_{4}[1,1]$ and the abelian part of $d_{4}[2]$ based on the direct QCD calculations  [^2] The corresponding values of $d_{4}[0,0,1],d_{4}[1,1]$ presented in Cvetic:2016rot , and Kataev:2022iqf , Goriachuk:2021ayq (Table 2 there) do not agree with the lhs in (3), the abelian part of their $d_{4}[2]$ does not agree with (10) with this $s_{4}$..
(iii) In the next step we will consider the contribution to $C_{F}\left(C_{F}^{2}x\right)$ that is known from Baikov:2010je , see the lhs of (11),
[…]
while its rhs is fulfilled by the weighted sum. Of course, the values of elements $d_{4}[.]$ in Eq.(8) satisfy this equality. Assuming that the value $d_{4}[0,0,1]$ is already confirmed from the condition (3) in item (i), one obtains an independent test for the partial sum $\hat{P}_{C_{F}^{3}x}\!\left(\beta_{1}d_{4}[0,1]+\,\beta_{0}d_{4}[1]\right)$ of the remaining elements,
[…]
To complete the confirmation for the 5-loop $d_{4}[.]$ elements, we need at least 3 independently calculated elements, e.g., $d_{4}[0,0,1],\,d_{4}[1,1]$ and, say, $d_{4}[0,1]$. In other words, taking into consideration the Casimir content of these elements, this parameterization consist of 4 independent scalar parameters333a similar parameterization was considered by A.Kataev, see A. Kataev’s talk at the session of RAS&JINR on 03.04.2024, https://indico.jinr.ru/event/4174/contributions/25722/. So to obtain the important element $d_{4}[0](c_{4}[0])$, one should take $d_{4}(x_{0})$ or $c_{4}(x_{0})$, all the powers of $\beta_{0}$ cancel in the corresponding $\{\beta\}$-expansion, see Eq.(2c,5b), and one arrives at
$d_{4}[0]$ $=$ $d_{4}(x_{0})-\beta_{2}(x_{0})\,d_{4}[0,0,1]-\beta_{1}(x_{0})\,d_{4}[0,1]\,,$ (12a)
$c_{4}[0]$ $=$ $c_{4}(x_{0})-\beta_{2}(x_{0})\,c_{4}[0,0,1]-\beta_{1}(x_{0})\,c_{4}[0,1]\,\,.$ (12b)
It is important if the elements $d_{4}[0,0,1],d_{4}[0,1]$ are already known, then the counter part elements $c_{4}[0,0,1],c_{4}[0,1]$ are also known from the corollaries Baikov:2022zvq of the CBK relation Baikov:2022zvq ; Chetyrkin:2022fqk like Eq.(7).
Let us mention here a rather rough test for elements $d_{3,4}[0]$ that is related to a special choice of Casimirs $C_{A},C_{F}$, see (25). The coefficients of perturbative expansion based on $SU_{c}(N)$ dynamics are the polynomials over the $C_{A},C_{F}$, let us consider these Casimirs as variables, which are related by the condition $07C_{A}+11C_{F}=$ 0 (that formally corresponds to the value $N=\sqrt{11}/5$). This linear condition, see Eqs.(5b, 5d), leads to the evident nullification of the coefficients $\beta_{0},\beta_{1}$,
$\beta_{0}(x_{0})=\beta_{1}(x_{0})=\beta_{0}(x_{1})=\beta_{1}(x_{1})=0\,.$ (13)
This condition radically simplifies the color structure of elements and in virtue of (13) leads
$\!\!\!\!\!\text{at}~{}7C_{A}+11C_{F}=0:$
[…]
The Eqs.(4c) for $d_{3}[0]$, and (8c) for $d_{4}[0,0,1]$, (8g) for $d_{4}[0]$ as well as the corresponding results in Goriachuk:2021ayq satisfy both the tests in Eqs.(14,15).
4 Reduction of $\{\beta\}$-expansion to a generalized NNA
Here we consider a conjecture about the hierarchy of contributions to the perturbation coefficients $d_{n},\,c_{n}$ based on the $\{\beta\}$–expansion, which generalizes the well-known NNA Broadhurst:1994se ; Beneke:1994qe . It also generalizes the results of the $\beta_{0}^{n}$-expansion Lovett-Turner:1995zwc and the $(\beta_{0}^{n},\beta_{1}\beta_{0}^{n-2})$ expansion in Ball:1995ni . This hierarchy rearranges the elements into a group that can provide the main contribution to the coefficients of the expansion $d_{n}(c_{n})$. It was first introduced in Mikhailov:2004iq (Sec.4) and was discussed in Bakulev:2010gm ; Kataev:2014jba . We will discuss the relations between the main term of this hierarchy, the topological structure of the diagrams corresponding to it in the abelian limit, and the conditions on the indices of the involved elements of the $\{\beta\}$-expansion.
4.1 Generalized NNA, illustration of $\beta_{0}$-dominance
Here we will use the ordering of contributions in powers of $\beta_{0}$ under the condition of “large $\beta_{0}$” (in QCD $\beta_{0}\sim 10$). More precisely, the common factor $\beta_{0}^{n-1}$ in any order $n$ will be taken out of the expansion coefficient $d_{n}$ to order all the elements over the inverse powers of $1/(\beta_{0})^{i}$ Mikhailov:2004iq . Recall that the well-known BLM Brodsky:1982gc trick that was a predecessor of the $\{\beta\}$-expansion, assumes for its efficiency the condition $\mid\beta_{0}d_{2}[1]\mid\gtrsim d_{2}[0]$, see the corresponding ratios in Eq.(16).
Then, we follow the “practical” observation $\beta_{i}=O(\beta_{0}^{i+1})$ that works at $n_{f}=0\div 5$, $i=1\div 4$, and invent the notation $b_{i}\stackrel{{\scriptstyle def}}{{=}}\beta_{i}/\beta_{0}^{i+1}=O(1)$, see (27). Of course, the last equality should be broken down at higher orders $i$ of pQCD due to the further factorial growth of $\beta_{i}$. So joining both these conditions, we take into account the terms with the coefficients $b_{i}$ together with others in the same order of $1/(\beta_{0})^{i}$, see the first brackets in Eqs.(17) and the first and the second brackets in Eqs.(18) below. We call this kind of rearrangement in the groups of the same $1/\beta_{0}$-order in brackets the generalized NNA (gNNA). The expansion coefficients of the $D_{A}$-function and BpSR $S^{Bjp}$ are presented below as the sums of these groups accompanied by their numerical estimates in inscriptions (at $n_{f}=3$, NNA estimates are shown from below in red),
$d_{2}=\beta_{0}\Big{[}\underbrace{d_{2}[1]}_{{\color[rgb]{1,0,0}\bm{2.77}}}+\overbrace{\frac{1}{\beta_{0}}d_{2}[0]}^{\bm{0.15}}\Big{]}\approx\beta_{0}[\bm{2.92}];~{}c_{2}=\beta_{0}\Big{[}\underbrace{c_{2}[1]}_{{\color[rgb]{1,0,0}\bm{-8}}}+\overbrace{\frac{1}{\beta_{0}}c_{2}[0]}^{\bm{1.63}}\Big{]}\approx\beta_{0}[\bm{-6.4}]$ (16)
[…]
$c_{3}$ $=$ $\beta_{0}^{2}\,\bigg{[}\overbrace{\Big{(}\underbrace{c_{3}[2]}_{{\color[rgb]{1,0,0}\bm{-25.6}}}+b_{1}\,c_{3}[0,1]\Big{)}}^{\bm{-25.9}}+\overbrace{\frac{1}{\beta_{0}}c_{3}[1]}^{-17.8}+\overbrace{\frac{1}{\beta_{0}^{2}}c_{3}[0]}^{27.7}\bigg{]}\approx\beta_{0}^{2}\,\left[\,\bm{-16}\,\right]$ (17b)
$\!\!\!\!\!\!d_{4}$ =\!= $\!\beta_{0}^{3}\!\bigg{[}\overbrace{\Big{(}\underbrace{d_{4}[3]}_{{\color[rgb]{1,0,0}\bm{8.7}}}+b_{1}\,d_{4}[1,1]+b_{2}\,d_{4}[0,0,1]\Big{)}}^{\bm{24}}+\overbrace{\frac{1}{\beta_{0}}(d_{4}[2]+b_{1}d_{4}[0,1])}^{34.3}+\overbrace{\frac{1}{\beta_{0}^{2}}d_{4}[1]}^{-6.8}+\overbrace{\frac{1}{\beta_{0}^{3}}d_{4}[0]}^{-34.35}\bigg{]}$ (18a)
[…]
$\!\!\!\!\!\!c_{4}$ =\!= $\!\beta_{0}^{3}\!\bigg{[}\overbrace{\Big{(}\underbrace{c_{4}[3]}_{{\color[rgb]{1,0,0}\bm{-89.6}}}+b_{1}\,c_{4}[1,1]+b_{2}\,c_{4}[0,0,1]\Big{)}}^{\bm{-112}}+\overbrace{\frac{1}{\beta_{0}}(c_{4}[2]+b_{1}c_{4}[0,1])}^{-95.5}+\overbrace{\frac{1}{\beta_{0}^{2}}c_{4}[1]}^{136.5}+\overbrace{\frac{1}{\beta_{0}^{3}}c_{4}[0]}^{9.4}\bigg{]}$ (18b)
$\approx$ $\beta_{0}^{3}\left[\,\bm{-61.7}\,\right]$
As it is seen, the leading $(1/\beta_{0})^{0}$-groups numerically dominate in all of the sums in (17,18), while the subleading groups for the most part mutually cancel one another, compare the lhs and the rhs there. The evident manifistation of gNNA vs NNA can be seen in the leading group of (18a). The numerical estimates of each of the elements $d_{2,3,4}[.],~{}c_{2,3,4}[.]$ are shown in Appendix B, where the diagrammatically confirmed results are marked in red. In addition, just the result of this leading group has been independently confirmed here in Sec.3, see the rhs of Eq.(3), which was based on the calculations of the QCD diagrams presented in Fig.2. Therefore, the correspondence of gNNA and the total sum in (18a) does not even depend on the detailed results in Eqs.(8b,8c) obtained with the help of QCDe. On the other hand, the contributions of the suppressed in $1/\beta_{0}$ the so-called “scale invariant” terms $d_{3,4}[0]$ almost cancel with the other subleading groups. At the same time, these $d_{i}[0]~{}(c_{i}[0])$ do not correspond by themselves to the final sum in the rhs neither in value nor in sign  [^4] The authors of PMC, see DiGiustino:2023jiq and references therein, propagate another, important role of the “scale invariant” elements $d_{n}[0]$ within the PMC approach.. We recognize that our conjecture is based on observations of a few examples only, and the reason of the aforementioned mutual cancellations is not clear; nevertheless, it looks reasonable to formulate the conjecture for the next orders of pQCD.
4.2 Conjecture of $\beta_{0}$-dominance
In the case of 6 loop Baikov:2022zvq , the leading gNNA group is presented in the first line in Eq.(19c) below. Following our conjecture, one needs to calculate restricted types of the corresponding diagrams to estimate the complete result $d_{5}$ based on (19c),
$d_{5}$ $=$ $\!\beta_{0}^{4}\!\bigg{[}\Big{(}\overbrace{d_{5}[4]+b_{1}d_{5}[2,1]+b_{1}^{2}\,d_{5}[0,2]+b_{2}d_{5}[1,0,1]+b_{3}d_{5}[0,0,0,1]}^{p(5-1)}\Big{)}+$ (19c)
$\frac{1}{\beta_{0}}\Big{(}d_{5}[3]+b_{1}d_{5}[1,1]+b_{2}d_{5}[0,0,1]\Big{)}+\frac{1}{\beta_{0}^{2}}\Big{(}d_{5}[2]+b_{1}d_{5}[0,1]\Big{)}+$
$\frac{1}{\beta_{0}^{3}}d_{5}[1]+\frac{1}{\beta_{0}^{4}}d_{5}[0]\bigg{]}\,.$
The number of elements of the leading group in Eq.(19c) is equal to the number of partitions $p$ depending on perturbation order $n$, e.g. , here – $p(5-1)=\bm{5}$ OEIS , see the detailed discussion of this correspondence in Baikov:2022zvq , Sec.4. So taking a natural series for the argument of the function $p$, one obtains $p(1,2,3,\bm{4},5,6,7,\ldots)=$ $1,2,3,\bm{5},7,11,15,\ldots$ For the order $n$ of pQCD the leading gNNA group consists of the number of partitions $p(n-1)$ elements that can be traced from the dressing of only one gluon line of the diagram. The indices of the elements $d_{n}[j_{1},\ldots,j_{n-1}]$ of the leading group satisfy the evident equality $11j_{1}+2j_{2}+\ldots\!+\!(\!n\!-1)\!j_{n-1}=n-$ 1, while for the different subleading groups the conditions for the indices are $1\,j_{1}+2\,j_{2}+\ldots+(n-1)\,j_{n-1}=n-2,n-3,\ldots$, where the shift from the number $n-1$ in the rhs determines the order of $1/\beta_{0}$ suppression see, e.g. , the groups in (19c, 19c).
If the elements of the gNNA group in (19c) are already obtained, they can be easily checked following the topological types of “abelian” diagrams in Fig.3,
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: The types of “abelian” diagrams corresponding to gNNA in order $a_{s}^{5}$ with all possible interchanges of the inner loops. (a) $C_{\text{F}}\,x^{4}$ $\sim\beta_{0}^{4}$; $C_{\text{F}}\,x^{3}$: (b), (c), (d) $\Rightarrow\beta_{3},\,\beta_{0}\beta_{2},\,\beta_{0}^{2}\beta_{1}$;  $C_{\text{F}}^{2}\,x^{2}$: (e),(f) $\Rightarrow\beta_{1}\beta_{1},\beta_{3}$.
where the chain of $n$ bubbles in Fig.3(a) definitely corresponds to only $d_{n}[n-1]$. The other diagrams – $(b,c,d)$ contribute to different combinations of the $\beta$-coefficients admissible for the leading group that has a projection on $C_{\text{F}}\cdot x^{3}$, while the diagrams $(e,f)\sim C_{\text{F}}^{2}\cdot x^{2}$. So one has
sum of diag.(b,c,d) $=$ $P_{C_{\text{F}}x^{3}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{2}\beta_{0}\,d_{5}[1,0,1]+\beta_{1}\beta_{0}^{2}\,d_{5}[2,1]\right);$ (20)
sum of diag.(e,f) $=$ $P_{C_{\text{F}}^{2}x^{2}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{1}^{2}\,d_{5}[0,2]\right)\,.$ (21)
It is clear that the diagrams similar those in Figs.2, 3 for the leading gNNA group can also be easily identified for the twist 2 coefficient function $C_{Bjp}$ of the observable $S^{Bjp}$.
To conclude, we present the leading gNNA for any order $n\geqslant 4$ as a weighted sum of $p(n-1)$ elements,
$d_{n}\!\!$ $=$ $\!\!\beta_{0}^{n-1}\bigg{[}\Big{(}\!\overbrace{d_{n}[n\!-\!1]+b_{1}d_{n}[n\!-\!3,1]+\!\ldots\!+b_{(n-2)}d_{n}[\underbrace{0,\ldots,1\!}_{n-1}]}^{p(n-1)}\!\Big{)}+\ldots+\frac{d_{n}[0]}{\beta_{0}^{n-1}}\bigg{]}\,.$ (22)
This group corresponds to the diagrams with only one dressed gluon line like those presented in Fig.3. On the contrary, the subleading gNNA groups correspond to the diagrams with more than one number of (dressed) gluon lines. It should be noted that taking into account the first two terms in the leading group of (22) to improve the BLM was first proposed in Ball:1995ni .
Using another language based on the hierarchy of $n_{f}$ orders, the application of gNNA requires taking into account in addition to the leading powers of $n_{f}$, a subseries of subleading orders also originating from different sources.
5 Conclusion
We have confirmed our results for the elements of the $\{\beta\}$-expansion in Baikov:2022zvq ; Mikhailov:2004iq : in order of $O(a_{s}^{3})$ – completely; in order of $O(a_{s}^{4})$ we confirmed two elements, $d_{4}[0,0,1],d_{4}[1,1]$, and the abelian part of $d_{4}[2]$ from Baikov:2022zvq , and for both the cases – basing on the known calculation Ball:1995ni within the standard QCD only.
We proposed a conjecture of $\beta_{0}$-dominance in $d_{n}$. The weighted sum of such elements $d_{n}[j_{1},j_{2},\ldots,j_{n-1}]$ dominates, whose indices satisfy the condition: $1\!\!1j_{1}\!+\!2j_{2}\!+\ldots\!+\!(\!n\!-1)\!j_{n-1}=n-$ 1, their contributions are not suppressed by powers of $1/\beta_{0}$. This condition one to one corresponds to allowing for the topological class of diagrams with renormalization of only one gluon line. The conjecture, in a lucky case, can significantly reduce the number of diagrams considered and make the laborious process of estimating physical quantities in high loops of calculations easier.
Acknowledgements.
The author is thankful to A. L. Kataev for drawing my attention to ref. Ball:1995ni and also for the constructive criticism of the paper, to N. Volchanskiy for the fruitful discussions. I am very grateful to K. Chetyrkin for the clarifying the important items of this consideration.
Appendix A QCD $\beta$-function
The well-known renormalization group equation for the QCD coupling constant $a_{s}(\mu^{2})\!=\!\frac{\alpha_{s}(\mu^{2})}{4\pi}$ reads
$\mu^{2}\frac{da_{s}(\mu^{2})}{d\mu^{2}}=-\beta(a)=-a_{s}^{2}(\mu^{2})\sum_{i\geq 0}\beta_{i}~{}a_{s}^{i}(\mu^{2})\,.$ (23)
The explicit expressions for the first coefficients of the $\beta$-function expansion with the underlined abelian terms read
$\beta_{0}$ $=$ $-\frac{4}{3}\,\underline{T_{R}n_{f}}+\frac{11}{3}\,C_{\text{A}}\,;\qquad$ (24a)
$\beta_{1}$ $=$ $-4\,\underline{C_{\text{F}}T_{R}n_{f}}+\frac{34}{3}\,C_{\text{A}}^{2}-\frac{20}{3}\,C_{\text{A}}T_{R}n_{f};$ (24b)
$\beta_{2}$ $=$ $\frac{44}{9}\underline{C_{F}(T_{R}n_{f})^{2}}+2\,\underline{C_{F}^{2}T_{R}n_{f}}+\frac{2857}{54}C_{A}^{3}-\frac{205}{9}C_{F}C_{A}T_{R}n_{f}-\frac{1415}{27}C_{A}^{2}T_{R}n_{f}$ (24c)
$+\frac{158}{27}C_{A}(T_{R}n_{f})^{2}\,;$
$\beta_{3}$ $=$ $\frac{1232}{243}\,\underline{C_{F}(T_{R}n_{f})^{3}}+\underline{(C_{F}T_{R}n_{f})^{2}}\left(\frac{1352}{27}-\frac{704}{9}\zeta_{3}\right)+46\underline{(C_{F})^{3}T_{R}n_{f}}+C_{A}^{3}T_{R}n_{f}\left(-\frac{39143}{81}+\frac{136}{3}\zeta_{3}\right)$ (24d)
$+C_{A}C_{F}T_{R}^{2}n_{f}^{2}\left(\frac{17152}{243}+\frac{448}{9}\zeta_{3}\right)+C_{A}C_{F}^{2}T_{R}n_{f}\left(-\frac{4204}{27}+\frac{352}{9}\zeta_{3}\right)+\frac{424}{243}C_{A}T_{R}^{3}n_{f}^{3}$
$+C_{A}^{2}C_{F}T_{R}n_{f}\left(\frac{7073}{243}-\frac{656}{9}\zeta_{3}\right)+C_{A}^{2}T_{R}^{2}n_{f}^{2}\left(\frac{7930}{81}+\frac{224}{9}\zeta_{3}\right)+C_{A}^{4}\left(\frac{150653}{486}-\frac{44}{9}\zeta_{3}\right)$
$+n_{f}\frac{d_{F}^{abcd}d_{A}^{abcd}}{N_{A}}\left(\frac{512}{9}-\frac{1664}{3}\zeta_{3}\right)+n_{f}^{2}\frac{d_{F}^{abcd}d_{F}^{abcd}}{N_{A}}\left(-\frac{704}{9}+\frac{512}{3}\zeta_{3}\right)$
$+\frac{d_{A}^{abcd}d_{A}^{abcd}}{N_{A}}\left(-\frac{80}{9}+\frac{704}{3}\zeta_{3}\right),$
with the $SU_{c}(N)$-group fundamental fermion invariants
$T_{R}=\frac{1}{2}\,;\;C_{\text{F}}=T_{R}\frac{N^{2}-1}{N}\,;\;C_{A}=N\,;\;N_{A}=2C_{\text{F}}C_{\text{A}}\equiv N^{2}-1\,;$ (25)
$d^{abc}d^{abc}=\frac{(N^{2}-4)}{N}N_{A}\,;~{}d_{F}^{abcd}d_{A}^{abcd}=\frac{N(N^{2}+6)}{48}N_{A}\,;$
$d_{F}^{abcd}d_{F}^{abcd}=\frac{N^{4}-6N^{2}+18}{96N^{2}}N_{A}\,;~{}d_{A}^{abcd}d_{A}^{abcd}=\frac{N^{2}(N^{2}+36)}{24}N_{A}\,,$ (26)
$d_{R}$ is the dimension of the quark color representation, $d_{R}=3$ in QCD and $n_{f}$ denotes the number of active flavors. At $n_{f}=3$,
$\beta_{0}=9;~{}~{}b_{1}\approx 0.79;~{}~{}b_{2}\approx 0.88;~{}~{}b_{3}\approx 1.9.$ (27)
Appendix B Numerical estimates of the elements $d_{2,3,4}[.],c_{2,3,4}[.]$ of $\{\beta\}$-expansion
We present the numerical estimates for the each of elements, discussed in Sec.4 for reader’s convenience. For the expressions $d_{2,3}[.]$ in Eq.(3,4) and similar ones $c_{2,3}[.]$ Baikov:2022zvq we have Kataev:2014zha ; Kataev:2014zwa ,
$\!\!\!\!{\color[rgb]{1,0,0}d_{2}[1]}$$\approx$ ${\color[rgb]{1,0,0}\bm{2.77}},~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}d_{2}[0]}={\color[rgb]{1,0,0}\bm{4/3}},~{}~{}~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}c_{2}[1]}={\color[rgb]{1,0,0}\bm{-8}},~{}~{}~{}~{}~{}~{}{\color[rgb]{1,0,0}c_{2}[0]}={\color[rgb]{1,0,0}\bm{44/3}};$ (28a)
$\!\!\!\!{\color[rgb]{1,0,0}d_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{12.41}},~{}~{}{\color[rgb]{1,0,0}d_{3}[0,1]}\approx{\color[rgb]{1,0,0}\bm{-4.799}},d_{3}[1]\approx 293.55,~{}~{}d_{3}[0]\approx-2295.84\,,$ (28b)
$\!\!\!\!{\color[rgb]{1,0,0}c_{3}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{-25.56}},~{}{\color[rgb]{1,0,0}c_{3}[0,1]\approx\bm{-0.43}},c_{3}[1]\approx-159.84,~{}c_{3}[0]\approx 2242.51\,.$ (28c)
From the expressions for $d_{4}[.]$ in Eq.(8) one has
${\color[rgb]{1,0,0}d_{4}[3]}$$\approx$ ${\color[rgb]{1,0,0}\bm{8.720}},~{}{\color[rgb]{1,0,0}d_{4}[1,1]}\approx{\color[rgb]{1,0,0}\bm{58.487}},~{}{\color[rgb]{1,0,0}d_{4}[0,0,1]}\approx{\color[rgb]{1,0,0}\bm{-34.955}},$ (29a)
${\color[rgb]{1,0,0}d_{4}[2]}$$\approx$ ${\color[rgb]{1,0,0}\bm{154.47}},~{}d_{4}[0,1]\approx 195.46,~{}d_{4}[1]\approx-550.54,$ (29b)
$d_{4}[0]$$\approx$ $-25105.61+20.5427\,n_{f}\approx-25043.97\,(\text{at}~{}n_{f}=3)$ (29c)
${\color[rgb]{1,0,0}c_{4}[3]}$$\approx$ ${\color[rgb]{1,0,0}\bm{-89.63}},~{}c_{4}[1,1]\approx-61.647,~{}{\color[rgb]{1,0,0}c_{4}[0,0,1]}\approx{\color[rgb]{1,0,0}\bm{29.722}},$ (30a)
$c_{4}[2]$$\approx$ $-707.165,~{}c_{4}[0,1]\approx-193.028,~{}c_{4}[1]\approx 11053.7,$ (30b)
$c_{4}[0]$$\approx$ $6871.02\,(\text{at}~{}n_{f}=3)$ (30c)
The elements marked in red in (28,29,30) have been confirmed, at least in their abelian part, by the diagrammatical calculations in the standard QCD in Secs.2, 3. The widely debated “zero” element $d_{4}[0]$ reads $-98.069+0.0802\,n_{f}$ in the case of the using the coupling constant $\alpha_{s}/\pi$ instead of $a_{s}$.